In a random sample of 300 patients, 21 experienced nausea. A drug manufacturer claims that fewer than 10% of patients who take its new drug for treating Alzheimer’s disease will experience nausea. Test this claim at the 0.05 significance level. What is the appropriate null hypothesis for this situation

1
Expert's answer
2022-05-31T15:32:53-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\ge0.10"

"H_a:p<0.10"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

Evidence:

Based on the information provided, the significance level is "\alpha = 0.05

," and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this left-tailed test is "R = \{z: z < -1.6449\}."

The z-statistic is computed as follows:



"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{21/300-0.1}{\sqrt{\dfrac{0.1(1-0.1)}{300}}}=-1.732"

Since it is observed that "z = -1.732<-1.6449= z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z<-1.732)= 0.041637," and since "p=0.041637<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p" is less than 0.10, at the "\alpha = 0.05" significance level.